How do you do y=-5x+12
y=1x-18

Answer by Theo(13342)

you have 2 equations.
y = -5x + 12
y = 1x - 18
subtract the first equation from the second equation to get:
0 = 6x - 30
add 30 to both sides to get 6x = 30
solve for x to get x = 5
that's your common solution to both equations.

y = -5x + 12 becomes y = -5(5) + 12 which becomes y = -25 + 12 which becomes y = -13.

y = x - 18 becomes y = 5 - 18 which becomes y = -13.

both equations give you y = -13 when x = 5.

you could also have simply set both equations equal to each other.

since y = -5x + 12 and y = x - 18, then by substution, you get -5x + 12 = x - 18

add 5x to both sides of this equation and add 18 to both sides of this equation to get 30 = 6x
solve for x to get x = 5.

same answer either way.